OpenStudy (anonymous):

What is a simpler form of the radical expression? 4^square root 81x^20 y^8

OpenStudy (anonymous):

\[\sqrt[4]{81x^{20}y^8}\] like that ?

OpenStudy (anonymous):

\[\sqrt[4]{81x ^{20}y ^{8}}\]

OpenStudy (anonymous):

yeah haha you're faster than me

OpenStudy (anonymous):

this is a lot easier than you think probably how many times does 4 go in to 20?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

right and so \[\sqrt[4]{x^{20}}=x^5\]

OpenStudy (anonymous):

using the same reasoning, what do you think \[\sqrt[4]{y^8}\]is ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

well actually \(y^2\) but you get the idea

OpenStudy (anonymous):

and all that is left to do is find \(\sqrt[4]{81}\) make a guess

OpenStudy (anonymous):

hint, it has to be small, since when you multiply it by itself 4 times you get 81

OpenStudy (anonymous):

9

OpenStudy (anonymous):

nope too big it is true that \(9^2=81\) but \(9^4=81\times 81\) go smaller

OpenStudy (anonymous):

3

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[3^4=3\times 3\times 3\times 3=9\times 9=81\]

OpenStudy (anonymous):

3[x^5]y^2

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

final answer: \[\sqrt[4]{81x^{20}y^8}=3x^5y^2\]the radical is gone

OpenStudy (anonymous):

yes, you don't need the parentheses though, just write them next to each other

OpenStudy (anonymous):

Thank you very much!

OpenStudy (anonymous):

yw wasn't even that hard was it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Sorry computer froze again, thank you again @satellite73